OpenStudy (anonymous):

help me solve this please .! .. The number of traffic accidents that occur on a particular stretch of road during a month follows a Poisson distribution with a mean of 8.9. Find the probability that less than three accidents will occur next month on this stretch of road. Answer choices..>>> 0.022777 0.977223 0.006752 0.993248

OpenStudy (tkhunny):

You should calculate that. p(0) + p(1) + p(2)

OpenStudy (anonymous):

?? im confused i have no idea what im doing

OpenStudy (tkhunny):

For the Poisson Distribution, \(p(n) = \dfrac{e^{-\lambda}\cdot \lambda^{n}}{n!}\). Does that look familiar?

OpenStudy (anonymous):

very familiar, but im not sure how to plug in

OpenStudy (tkhunny):

Simply evaluate p(0), p(1), and p(2). You'll be done after you add them up.

OpenStudy (tkhunny):

\(\lambda=8.9\) -- Just a property of this Poisson Distribution, since you are given the Mean.

OpenStudy (anonymous):

hmm ok i got .90 something so im thinking its either B or D>?

OpenStudy (tkhunny):

e^{-8.9} = 0.000136388926482 \(\dfrac{\lambda^{0}}{0!} = 1\) \(\dfrac{\lambda^{1}}{1!} = 8.9\) \(\dfrac{\lambda^{2}}{2!} = 39.605\) 0.000136388926482*(1 + 8.9 + 39.605) = 0.00675 How did you manage ".90 something"? Frankly, the mean is 8.9. I sincerely hope we don't get 0.90 in the first three probabilities.